We wish to estimate the mean serum indirect bilirubin level of 4-day-old infants. The mean for a sample of 16 infants was found to be 5.98 mg100 cc. Assume that bilirubin levels in 4-day-old infants are approximately normally distributed with a standard deviation of 3.5 mg100 cc.


1
Expert's answer
2021-06-07T14:39:37-0400

"n=16 \\

\bar{x}= 5.98 \\

s=3.5 \\

H_0: \mu= 4 \\

H_1: \mu > 4"

Test statistic:

"t = \frac{\bar{x}- \mu}{s / \sqrt{n}} \\

= \frac{5.98-4}{3.5 / \sqrt{16} } \\

= 2.2629"

p-value = 0.01945

p-value<0.05

Reject H0.

It is concluded that, population mean is greater than 4.

The 95 % confidence interval for the population mean:

"Z_{0.05/2}=1.96 \\

CI = \bar{x}±Z_{0.05/2} \times \frac{s}{\sqrt{n}} \\

= 5.98 ±1.96 \times \frac{3.5}{\sqrt{16}} \\

= 5.98 ± 1.715 \\

(4.265, 7.695)"